Basic Concepts

Roberts Company is considering an investment in equipment that is capable of producing more efficiently than the current technology. The outlay required is $2,200,000. The equipment is expected to last five years and will have no salvage value. The expected cash flows associated with the project are as follows:

Year Cash Revenues Cash Expenses
1 $2,960,000 $2,300,000
2 2,960,000 2,300,000
3 2,960,000 2,300,000
4 2,960,000 2,300,000
5 2,960,000 2,300,000

The present value tables provided in Exhibit 19B.1 and Exhibit 19B.2 must be used to solve the following problems.

Required:

1. Compute the project’s payback period. If required, round your answer to two decimal places.
years

2. Compute the project’s accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16" in the answer box).
%

3. Compute the project’s net present value, assuming a required rate of return of 10 percent. When required, round your answer to the nearest dollar.
$

4. Compute the project’s internal rate of return. Enter your answers as whole percentage values.

Between  % and  %
Answer #1

Solution 1:

Annual cash inflows = $2,960,000 - $2,300,000 = $660,000

Payback period = Initial investment / Annual cash inflows = $2,200,000 / $660,000 = 3.33 years

Solution 2:

Annual income = Annual cash inflows - Depreciation = $660,000 - ($2,200,000 / 5) = $220,000

Average investment = (Initial investment + Salvage value) / 2 = ($2,200,000 + 0) / 2 = $1,100,000

Accounting rate of return = Average annual income / Average investment - $220,000 / $1,100,000 = 20%

Solution 3:

Computation of NPV
Particulars Period Amount PV factor at 10% Present Value
Cash outflows:
Initial investment 0 $2,200,000.00 1 $2,200,000
Present Value of Cash outflows (A) $2,200,000
Cash Inflows
Annual cash inflows 1-5 $660,000.00 3.7910 $2,502,060
Present Value of Cash Inflows (B) $2,502,060
Net Present Value (NPV) (B-A) $302,060

Solution 4:

Present value factor at IRR = $2,200,000 / $660,000 = 3.333

Refer PV factor table, this factor falls between 15% and 16%

Therefore IRR lies between 15% and 16%.